OpenStudy (anonymous):

Hello! Need some help figuring out MC, TVC, AFC,and ATC. A firm producing100 units of output at a total cost of $800. The firm's average variable cost is $5 per unit.

OpenStudy (anonymous):

First, you need to know what they stand for. 1. MC = Marginal Cost 2. TVC = Total Variable Cost 3. AFC = Average Fixed Cost 4. ATC = Average Total Cost Question: A firm is producing 100 units of output at a total cost of $800. The firm's average variable cost is $5 per unit. 1. Marginal cost = the cost of producing 1 extra unit of output. Total cost = $800. Number of units = $100 Divide 800 by 100 = 8 it costs $8 on average to produce 1 unit. To produce an extra unit, it should cost another $8... So I'd say $8 is the MC 2. Total Variable Cost = total cost - fixed cost, although you have to change it up a bit for this questions and use application: The average variable cost = $5 per unit Total number of units produced = 100 units. Multiply 5 X 100 = 500 So TVC = $500 3. Average fixed Cost = the fixed cost per unit of output. Relying on these first two questions, total cost = $800. Total cost - total variable cost = total fixed cost 800 - 500 = 300 TFC = 300. However, we're finding the average (per unit) In total, there are 100 units. Therefore, the fixed cost for 1 unit of output = 300/100 = 3 AFC = $3 4. Average Total Cost = total cost / quantity produced. Total cost = $800 quantity produced = $100 800 / 100 = 8 so average total cost = $8 ~ I would recheck the MC, as I'm not entirely sure that I'm right there; however, I would not know what to do otherwise. Hope this helps!